OpenStudy (anonymous):

A random variable X has mean equal to 4/5 and variance equal to 2/75. Find approximations to the mean and variance of X^2?

OpenStudy (anonymous):

is the mean of X^2=16/25 +2/75 =2/3

OpenStudy (anonymous):

The X^2 refers to the second moment i think... not just X squared

OpenStudy (phi):

I interpret this question as finding the mean and variance of f(x)= x^2 where x is a random variable with a normal distribution (0.8, sqrt(2/75) ) (mean,std) If that is what it's asking, have you learned any way to approximate the variance of a random function? See http://en.wikipedia.org/wiki/Variance#Approximating_the_variance_of_a_function

OpenStudy (phi):

Do you know var(x) = E(x^2) - ( E(x) )^2 ? they give us var(x) and E(x) , so you can solve for E(x^2), which is the mean of the random function x^2

OpenStudy (anonymous):

This is the full question. Maybe it will make what is being asked a lot clearer

OpenStudy (anonymous):

Yh we need to use the approximation of random function method... just scrolled through my lectures and found it.